d) Assume that interest rate parity exists. You expect that the one-year nominal interest rate in the U.S. is 1.995%, while the one-year nominal interest rate in Australia is 3.695%. The spot rate of the Australian dollar is USD0.6939. You will need 15 million Australian Dollars in one year.

Today, you purchase a one-year forward contract in Australian Dollars. Estimate how many U.S. Dollars (USD) will you need in one year to fulfill your forward contract.
Answer #1

According to the interest rate parity,

Exchange rate after 1 year = Spot rate*(1+Interest rate in US)/(1+Interest rate in Australia)

Exchange rate after 1 year = 0.6939*1.01995/1.03695 = USD 0.6825

15 million Australian dollar equivalent in USD after 1 year = 15*0.6825 million USD = USD 10.2378 million

Hence you willl need 10.2378 million USD